OpenStudy (marcelie):

Help Please !!! Each side of a square is increasing at a rate of 6 cm/ s . at what rate is the area of the square increasing when the area of the square is 12 cm ^2?

OpenStudy (marcelie):

@Directrix

OpenStudy (dumbcow):

\[A = x^2\] Differentiate with respect to time using the chain rule \[\frac{dA}{dt} = \frac{dA}{dx} * \frac{dx}{dt}\] \[\frac{dA}{dt} = 2x (6) = 12x\] Finally when Area is 12, the side length (x) is sqrt(12) \[\frac{dA}{dt} = 12 \sqrt{12} = 24 \sqrt{3}\]

OpenStudy (marcelie):

okay then ?

OpenStudy (marcelie):

my answer sheet says 48 how did they get that

OpenStudy (dumbcow):

48 is wrong its 24sqrt(3)

OpenStudy (marcelie):

okay so then ty..

OpenStudy (marcelie):

@satellite73 plz help me im confuse about this ..